OpenStudy (anonymous):

Write an equation of a line parallel to y = 2 over 3x −5 in slope-intercept form that passes through the point (−6, −1). y= 2 over 3x + 3 y= 2 over 3x − 6 y= −3x − 1 y= 3x + 6

OpenStudy (aripotta):

parallel lines have the same slope. what is the slope of the given line?

OpenStudy (misty1212):

\[\huge y =\color{red}{\frac{2}{3}}x-5\]

OpenStudy (misty1212):

like that?

OpenStudy (anonymous):

Yes

OpenStudy (aripotta):

so what's the slope?

OpenStudy (misty1212):

hint: it is big and red !

OpenStudy (anonymous):

Im Not Sure.

OpenStudy (anonymous):

Oh Okay Lol Thanks!

OpenStudy (aripotta):

so we know our answer will have a slope of 2/3, and go through the point (-6, -1). we can plug that information into point-slope form, which is y - y1 = m(x - x1) where y1 is the y-coordinate of the point, x1 is the x-coordinate of the same point, and m is slope what do you get once you plug that in?

OpenStudy (anonymous):

Im Confused Of How To Plug It In

OpenStudy (anonymous):

@aripotta

OpenStudy (aripotta):

"plugging in" is just another word for substitute

OpenStudy (aripotta):

so for m, we'll "plug in" 2/3. get it?

OpenStudy (anonymous):

Yes I Know That But Im Confused Of The Order To Plugging It In.

OpenStudy (aripotta):

what's y1?

OpenStudy (aripotta):

our point is (-6, -1). y1 is the y part of that point. and x1 is the x part of that point.

OpenStudy (aripotta):

points are in the form of (x,y)...

OpenStudy (aripotta):

are you there? @jaygamble